OpenStudy (anonymous):

The average of two numbers is 18. One of the numbers is twice the other. What is the larger number?

OpenStudy (campbell_st):

let the 1st number = x so the other number is 2x then (x + 2x)/2 = 18 or (3x)/2= 18 solve for x

OpenStudy (anonymous):

Ok, so we have two numbers: \(a, b\) We know \(\Large \frac{a+b}{2} = 18\) and \(a = 2b\) This is a system of equations. Do you know how to solve for them?